OpenStudy (anonymous):

The table shows the outputs y for different inputs x: Input (x) 5 6 7 8 Output (y) 2 5 8 11 Part A: Do the data in this table represent a function? Justify your answer. (3 points) Part B: Compare the data in the table with the relation f(x) = 2x + 13. Which relation has a greater value when x = 7? (2 points) Part C: Using the relation in Part B, what is the value of x if f(x) = 75? (5 points)

OpenStudy (anonymous):

can somebody help me with part B

OpenStudy (triciaal):

you have a function when there is only one output value for a given input

OpenStudy (unklerhaukus):

The column corresponding to x =7 in the table, is here | Input | (x) 5 6 [7] 8 Output (y) 2 5 [8] 11 | | The value of the function: f(x) = 2x + 13 , evaluated at x=7 is f(7) = 2(7) + 13 =

OpenStudy (triciaal):

Which relation has a greater value when x = 7? as shown above which is greater 8 or 2(7) + 13 ?

OpenStudy (triciaal):

C) what is the value of x if f(x) = 75 f(x) = 2x + 13 = 75 solve for x

OpenStudy (triciaal):

do you know how to solve for x?